# Transductive Sample Complexities Are Compact

### v2

@@ 4 Conclusion
Julian Asilis was supported by the Simons Foundation and by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1842487. Siddartha Devic was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Shaddin Dughmi was supported by NSF Grant CCF-2009060. Vatsal Sharan was supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Shang-Hua Teng was supported in part by NSF Grant CCF-2308744 and the Simons Investigator Award from the Simons Foundation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any of the sponsors such as the NSF.
@@ Appendix A Proper metric spaces
Appendix B Omitted proofs
B.1 Proof of Theorem 3.8
@@ Appendix B Omitted proofs
Then, invoking Zorn’s lemma, $\mathcal{P}$ contains a maximal element $\phi_{\max}$. By Lemma B.2, it must be that $\phi_{\max}$ does not leave any function in $R$ unassigned, otherwise it could be augmented with an additional apportionment, contradicting maximality. Thus there exists a total assignment of $(\epsilon+\delta)$-apportionments which is finitely satisfiable, completing the argument. ∎
@@
The central observation is that defining a learner $\mathcal{A}$ for $\mathcal{H}$ which attains error $\leq 2\epsilon+\delta$ amounts precisely to assigning each variable in $L$ to a value in $\mathcal{Y}$ such that the functions in $R$ are all maintained below $2\epsilon+\delta$. By the premise of the theorem, this collection of functions and variables is finitely satisfiable. That is, for each finite $R^{\prime}\subseteq R$, there exists an assignment of all variables $L\to\mathcal{Y}$ such that all functions in $R^{\prime}$ are maintained below $\epsilon$ [^4] Strictly speaking, the definition of error rate involves an infimum, meaning we are only guaranteed that the functions in $R^{\prime}$ can be maintained arbitrarily close to $\epsilon$. It is straightforward to see that this suffices for our purposes, however, as Lemma B.1 results in the addition of an arbitrarily small term to $\epsilon$ anyway. Then, by Lemma B.1, there exists an assignment of $(\epsilon+\delta/3)$-apportionments to each function in $R$ which is finitely satisfiable.

### v3

@@ 4 Conclusion
Julian Asilis was supported by the Simons Foundation and by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1842487. Siddartha Devic was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Shaddin Dughmi was supported by NSF Grant CCF-2009060. Vatsal Sharan was supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Shang-Hua Teng was supported in part by the Simons Investigator Award from the Simons Foundation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any of the sponsors such as the NSF.
@@ Appendix A Proper metric spaces
Appendix B Distribution-family Learning
The analysis of PAC learning with respect to more flexible distribution classes than the realizable and agnostic cases falls largely under the purview of distribution-family learning [Benedek and Itai, 1991]. Formally, a problem in distribution-family learning of a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ is defined by a family of distributions $\mathbb{D}$ over $\mathcal{X}$, such that unlabeled datapoints are drawn from a distribution $D\in\mathbb{D}$ and labeled by a hypothesis $h\in\mathcal{H}$ (in the realizable case) or arbitrarily (in the agnostic case).
Notably, distribution-family learning has infamously resisted any characterization of learnability, combinatorial or otherwise, for the 40 years since its inception. In fact, there is some evidence to suggest that no such characterization may exist [Lechner and Ben-David, 2023]. Furthermore, it is a setting in which uniform convergence fails to characterize learning, rendering ineffective many of the standard and most celebrated techniques of learning theory.
Definition B.1.
Definition B.1 is sufficiently flexible that we may apply it to distribution-family learning with $\mathcal{Z}=\mathcal{X}$ or to PAC learning over arbitrary distribution classes $\mathbb{D}$ with $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. Though it may appear overly restrictive at first glance, note that well-behavedness is satisfied not only for ordinary PAC learning in the agnostic and realizable cases, but also for learning of partial concept classes in the realizable case [Alon et al., 2022] and for the EMX learning of Ben-David et al. [2019]. In particular, though EMX learning is not presented as a supervised learning problem in Ben-David et al. [2019], it can be seen as a binary classification problem over a domain $\mathcal{X}$ for which $\mathcal{H}$ consists of those functions outputting finitely many 1’s and $\mathbb{D}$ contains all realizable, discrete distributions placing all $\mathcal{Y}$-mass on the label 1.
Proposition B.2.
Proof.
The proof is nearly identical to that of [Asilis et al., 2024, Proposition 3.6]. In particular, let $m_{\operatorname{Exp},\mathcal{H}}$ denote the sample complexity of learning $\mathcal{H}$ in the expected error regime (i.e., $m_{\operatorname{Exp},\mathcal{H}}(\epsilon)$ equals the number of datapoints needed to incur expected error at most $\epsilon$). Then we have
The first inequality follows immediately from the fact that the loss function is bounded above by 1. The second inequality follows from a repetition argument, i.e., a learner attaining expected error $\leq\epsilon/2$ on samples of size $n$ can be boosted to attain expected error $\leq\epsilon$ with probability $\geq 1-\delta$ by using an additional factor of $O(\log(1/\delta))$ many samples, as described in [Daniely and Shalev-Shwartz, 2014].
The first inequality follows inequality from a standard leave-one-out argument of Haussler et al. [1994]. The second inequality follows from the fact that for any $S=((x_{1},y_{1}),\ldots,(x_{n},y_{n}))$, there exists an $m_{n}\in\mathbb{N}$ such that $m_{n}$ many independent draws from the uniform distribution over the entries of $S$ has probability at least $\frac{1}{\epsilon}$ of containing exactly $n-1$ elements of $S$, as detailed in [Asilis et al., 2024, Lemma A.1]. Crucially, the uniform distribution over $S$ lies in $\mathbb{D}$ owing to well-behavedness of $\mathbb{D}$. The claim follows from both of the established chains of inequalities. ∎
Note that Proposition B.2 holds for the natural definition of transductive learning with respect to $\mathbb{D}$, i.e., in which the adversary must select a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$ which lie in the support of some $D\in\mathbb{D}$. It is now immediate from the proof of Theorem 3.6 that distribution-family learning is a setting in which the transductive sample complexity of learning is $\mathcal{H}$ equals the sample complexity of learning its most challenging finite projections. In the following, we let $\mathbb{D}|_{X}\subseteq\mathbb{D}$ denote the distributions of $\mathbb{D}$ which place full measure on $X\subseteq\mathcal{X}$.
Theorem B.3.
  • 1.
  • 2.
As in Section 3.4, Proposition B.2 applied to Theorem B.3 immediately yields an almost-exact form of compactness for distribution-family PAC learning with realizable and well-behaved distribution classes. This demonstrates that the learnability of even problems as exotic as EMX learning can be detected by examining all their finite projections, provided that no restrictions are placed upon learners. In Ben-David et al. [2019], however, learners were required to only emit hypotheses in $\mathcal{H}$. Our work demonstrates that the nature of their undecidability result — in which the learnability of a class $\mathcal{H}$ is determined entirely by its cardinality, despite all its projections being easily learned — could otherwise not appear in supervised learning with metric losses.
Appendix C Omitted proofs
C.1 Proof of Theorem 3.8
@@ Appendix C Omitted proofs
Then, invoking Zorn’s lemma, $\mathcal{P}$ contains a maximal element $\phi_{\max}$. By Lemma C.2, it must be that $\phi_{\max}$ does not leave any function in $R$ unassigned, otherwise it could be augmented with an additional apportionment, contradicting maximality. Thus there exists a total assignment of $(\epsilon+\delta)$-apportionments which is finitely satisfiable, completing the argument. ∎
@@
The central observation is that defining a learner $\mathcal{A}$ for $\mathcal{H}$ which attains error $\leq 2\epsilon+\delta$ amounts precisely to assigning each variable in $L$ to a value in $\mathcal{Y}$ such that the functions in $R$ are all maintained below $2\epsilon+\delta$. By the premise of the theorem, this collection of functions and variables is finitely satisfiable. That is, for each finite $R^{\prime}\subseteq R$, there exists an assignment of all variables $L\to\mathcal{Y}$ such that all functions in $R^{\prime}$ are maintained below $\epsilon$ [^4] Strictly speaking, the definition of error rate involves an infimum, meaning we are only guaranteed that the functions in $R^{\prime}$ can be maintained arbitrarily close to $\epsilon$. It is straightforward to see that this suffices for our purposes, however, as Lemma C.1 results in the addition of an arbitrarily small term to $\epsilon$ anyway. Then, by Lemma C.1, there exists an assignment of $(\epsilon+\delta/3)$-apportionments to each function in $R$ which is finitely satisfiable.